OpenStudy (anonymous):

sum is -6 product is 8

OpenStudy (anonymous):

x + y=-6 xy=8 answer is 4 and 2

OpenStudy (anonymous):

sorry -4 and -2

OpenStudy (anonymous):

that's ok !!!.. thnk you

OpenStudy (anonymous):

so x+y=-6 xy=8 so y=8/x x+8/x=-6 so \[(x^2+8)/x = -6 \] then x^2+8=-6x x^2+6x+8=0 (x+4)(x+2)=0 x= -4, -2